A parent rewards a child with 50 cents for each correctly solved mathematics problem and fines the child 30 cents for each incorrectly solved problem. If the child nets $22.00 after 100 problems, how many problems were solved correctly?
a. 30
b. 35
c. 45
d. 53
e.

Answer by checkley75(3666)

50x-30(100-x)=2200
50x-3000+30x=2200
80x=2200+3000
80x=5200
x=5200/80
x=65 correct answers.
100-65=35 incorrect answers.
proof
50*65-30*35=2200
3250-1050=2200
2200=2200